### v2

# Study of Flat Spectrum Radio Quasars and BL Lacertae Objects as Sources of Diffusive Ultra High-Energy Cosmic Rays

Abstract
We examine whether Flat Spectrum Radio Quasars (FSRQs) and BL Lacertae objects (BL Lacs) can act as plausible astrophysical sources of diffuse ultra-high-energy cosmic rays (UHECRs). Using realistic luminosity-dependent density evolution (LDDE) functions derived from observed gamma-ray luminosity functions for FSRQs and BL Lacs, we calculate the redshift evolution of the cosmic ray source population through integrated luminosity functions. The diffuse UHECRs flux from these sources is modelled by propagating nuclei through extragalactic space, including energy losses from interactions with cosmic photon backgrounds. The resulting UHECRs spectra are compared with observational data from the Pierre Auger Observatory and the Telescope Array, with fluxes normalised at reference energies. We also construct illustrative full skymaps of the integral UHECRs flux from $10$ EeV to $100$ EeV using the HEALPix framework, assuming a uniform distribution of synthetic sources and including energy-dependent magnetic diffusion. The regions of higher flux in the maps arise naturally from propagation effects and the limited UHECRs horizon. Our results indicate that LDDE-modelled AGNs could contribute significantly to the observed diffuse UHECRs flux and provide constraints on their role as potential sources.
Ultra-High Energy Cosmic Rays; Flux; Flat Spectrum Radio Quasars (FSRQs); BL Lacertae
I Introduction
The origin of ultra high-energy cosmic rays (UHECRs; $E\gtrsim 10^{17}$ eV) remains as one of the outstanding puzzles in modern astrophysics (for reviews, see [1, 2]). Unlike photons and neutrinos, the arrival directions of UHECRs are obscured due to deflections in galactic and extragalactic magnetic fields [3, 4], complicating the task of identifying their sources. Although the observed features in the UHECRs spectrum suggest an extragalactic origin, the precise energy at which the transition between galactic and extragalactic contributions occurs is still uncertain (e.g., see [5]). The Pierre Auger Observatory (Auger) in Argentina [6] and the Telescope Array (TA) in Utah, USA [7], currently provide the most detailed measurements of these particles.
A few spectacular events recorded in different UHECR experiments already highlight the involvement of extreme energies in such events. E.g., the “Oh-My-God” particle, detected by the Fly’s Eye experiment, carried an estimated energy of $\approx 320$ EeV ($1$ EeV $10^{18}$ eV), making it the most energetic event recorded via the fluorescence technique [8]. More recently, the TA Collaboration reported the “Amaterasu” event on May 27, 2021, which is the most energetic cosmic ray (CR) ever detected with a surface detector array. Its reconstructed energy of $244\pm 29\text{\ (stat)}^{+51}_{-76}\text{\ (sys)}$ EeV and arrival direction (RA, Dec) of ($255.9^{\circ}\pm 0.6^{\circ}$, $16.1^{\circ}\pm 0.5^{\circ}$) places it among the most energetic particles observed to date [9]. Such extreme-energy events raise the fundamental question about their acceleration mechanisms and astrophysical origins.
Propagation effects, i.e., the effects of the environmental situations on CRs while they propagate through space, also play a crucial role in their properties, including flux, energy and arrival directions from their sources. At the highest energies, UHECRs are expected to suffer significant attenuation due to the Greisen–Zatsepin–Kuz’min (GZK) suppression [10, 11], arising from photopion production with the cosmic microwave background (CMB) photons. The attenuation length of protons at these energies is of order $\sim 10$ Mpc (e.g., see [12]). Composition studies add further complexity: while TA results suggest predominantly light primaries at $E\gtrsim 10^{19}$ eV, significant uncertainties still remain [13].
Several astrophysical and cosmological scenarios have been proposed to account for the sources of UHECRs, which include ultraheavy nuclei [14], binary neutron star mergers [15], the decay of superheavy dark matter [16], and transient events in unresolved galaxies [17]. In addition, cosmological scenarios based on modified or alternative gravity theories provide an alternative framework to explain the acceleration, propagation, and large-scale distribution of UHECRs in the evolving Universe [18, 19]. Moreover, UHECRs are powerful probes of new physics such as Lorentz invariance violation (LIV) [20], and this possibility has also been explored for the Amaterasu event [21]. Recent high-precision measurements of UHECRs reveal several characteristic features: a hardening near the ankle at $\sim 5$ EeV [22, 23, 24], a subsequent steepening at $\sim 13$ EeV [23, 25], and a strong suppression beyond $\sim 50$ EeV [26, 27]. Composition studies based on the depth of shower maximum, $X_{\rm max}$, show that UHECRs are relatively light at a few EeV and become progressively heavier with increasing energy. Moreover, the narrow dispersion in $X_{\rm max}$ above the ankle suggests that the mass distribution at a given energy is relatively tight.
When these observations are interpreted within astrophysical source models, they indicate that two distinct source populations are required to explain the UHECRs flux above $\sim 1$ EeV and across the ankle [28, 29, 30]. The low-energy (LE) population, dominant below a few EeV, is likely composed of light and intermediate nuclei (H, He, N) with steep spectra $\propto E^{-\gamma_{\rm L}}$, where $\gamma_{\rm L}\simeq 3-3.7$, and with an only loosely constrained rigidity cutoff. At higher energies, the high-energy (HE) population becomes dominant, consisting of progressively heavier nuclei. This component is expected to have a much harder injection spectrum $\propto E^{-\gamma_{\rm H}}$, with $\gamma_{\rm H}<1$, and in some scenarios even negative spectral indices. Together, these populations provide a consistent description of the observed UHECRs spectrum and composition, linking spectral features to the physics of acceleration and propagation.
The study of radio-loud (RL) active galactic nuclei (AGN), particularly blazars with strongly relativistic and beamed jets, provides a unique method to probe jet activity, black hole spin, and the role of major mergers in galaxy evolution. A key statistical tool in this context is the luminosity function (LF) of blazars, defined as the number of blazars per comoving volume within a given luminosity interval, and its evolution with redshift. The Fermi Gamma-ray Space Telescope (Fermi-LAT) has revolutionized this field: thanks to its sensitivity and all-sky uniform coverage, it has detected hundreds of blazars spanning from the local Universe up to $z=3.1$ [31]. The blazar LF not only constrains their contribution to the diffuse extragalactic $\gamma$-ray background but also helps establish connections between blazars and their parent AGN populations [32, 33]. Extensive multiwavelength studies of blazars have been carried out in the radio [34, 35] and soft X-ray [36, 37, 38, 39, 40, 41] bands. Flat-spectrum radio quasars (FSRQs), the most luminous subclass of blazars, exhibit strong positive cosmological evolution, meaning they were more numerous in the past [34]. This trend persists up to a luminosity-dependent redshift cutoff (e.g., see [41, 42]). Interestingly, FSRQs follow an evolutionary pattern similar to that of X-ray-selected, radio-quiet AGNs [43, 44, 45], suggesting possible common mechanisms driving their cosmic evolution.
Another important sub-population of blazars are BL Lacertae (BL Lac) objects (e.g., see [46]). They represent an extreme class of AGNs characterized by highly variable emission, most likely produced by relativistic jets aligned very close to our line of sight. BL Lac objects are distinguished from their FSRQ counterparts by their optical spectra, which lack significant emission lines with equivalent width $5\AA$ (e.g., see [47, 48]). Their spectra are typically dominated by a power-law continuum, indicating either exceptionally strong non-thermal jet emission (aligned nearly directly with our line of sight), or unusually weak thermal disk and broad-line emission, the latter plausibly linked to low accretion activity [49].
Considering the said interesting nature of emission with high redshift values of FSRQs and BL Lac objects, we aim to investigate these objects as plausible sources of UHECRs from the perspective of the characteristics of UHE fluxes from them. Thus, the remainder of the paper is organised as follows. Section II is divided into two subsections. In Subsection II.1, we describe the evolution of FSRQ sources, while Subsection II.2 focuses on the evolution of BL Lac objects. Section III presents the basic equations governing the propagation of UHECRs. In Section IV, we provide the flux calculation equations, discuss the corresponding mixed composition results, generate skymaps, identify hotspots, and locate nearby galaxies within $1^{\circ}$ angular separation. Finally, Section V summarises the findings and draws conclusions of the study.
II Evolution of AGN Sources
In this section, we describe the evolution of two classes of AGN sources, viz., the FSRQs and BL Lac objects, which we consider as potential sources of UHECRs, as mentioned already. First, we describe the FSRQs’ evolution, and in the next the evolution of the BL Lac objects is discussed as follows.
II.1 Flat Spectrum Radio Quasars (FSRQs)
The space density of radio-quiet AGNs is observed to reach a maximum at intermediate redshifts. The epoch of this ‘redshift peak’ is correlated with the luminosity of the sources (e.g., see [43, 44]). This peak likely reflects the combined effects of supermassive black hole (SMBH) growth over cosmic time and a decline in fuelling activity as the frequency of major mergers decreases at later times. To investigate whether similar behaviour occurs in the Fermi-LAT FSRQ population, Ajello et al. fit the data using a pure luminosity evolution (PLE) model as described in Ref. [50] and is given as
$\Phi(L_{\gamma},z)=\Phi(L_{\gamma}/e(z)),$ (1)
where
$\Phi(L_{\gamma}/e(z=0))=\frac{dN}{dL_{\gamma}}=\;$ $\frac{A}{2.303\,L_{\gamma}}$
$\times\left[\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma_{1}}+\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma 2}\right]^{-1}$ (2)
and
$e(z)=(1+z)^{k}e^{z/\xi}.$ (3)
In the PLE model, the comoving luminosity function $\Phi(L_{\gamma},z)$ is assumed to evolve only in luminosity, such that the redshift dependence enters through the evolution factor $e(z)$. Here, $L_{\gamma}$ denotes the $\gamma$-ray luminosity of the source and $\Phi(L_{\gamma},z)$ gives the number density of sources per unit luminosity. It is to be noted that the observed $\gamma$-ray luminosities of blazars are affected by Doppler boosting and hence $L_{\gamma}$ is related with the intrinsic luminosity $L_{\gamma}^{\prime}$ as $L_{\gamma}=(\delta_{\rm e}^{6}/\Gamma_{\rm e}^{2})L_{\gamma}^{\prime}$ [50, 51], where $\delta_{\rm e}$ is the Doppler factor of the jet, which quantifies the enhancement of the observed luminosity $L_{\gamma}$ due to relativistic motion, and $\Gamma_{\rm e}$ is the Lorentz factor. The local luminosity function at $z=0$ is characterized by a broken power-law with a normalization factor $A$, break luminosity $L_{*}$, low-luminosity slope $\gamma_{1}$, and high-luminosity slope $\gamma_{2}$, where $\gamma_{1}$ controls the behavior of the faint-end population ($L_{\gamma}\ll L_{*}$) and $\gamma_{2}$ determines the fall-off at the bright end ($L_{\gamma}\gg L_{*}$). The luminosity evolution factor, $e(z)=(1+z)^{k}\exp(z/\xi)$, describes how the characteristic luminosity shifts with redshift, where $k$ sets the strength of luminosity evolution and $\xi$ controls the redshift scale at which the evolution transits or saturates. As a simple PLE function model does not adequately reproduce the Fermi-LAT data, and given evidence for a luminosity-dependent evolution of the redshift peak, we fit the Fermi-LAT FSRQ sample using a luminosity-dependent density evolution (LDDE) model. In this approach, the evolution occurs mainly in density, with the redshift peak depending on luminosity. The LDDE model is parametrised as in Ref. [50], which is
$\Phi(L_{\gamma},z)=\Phi(L_{\gamma})\times e(z,L_{\gamma}),$ (4)
where
$e(z,L_{\gamma})=\left[\left(\frac{1+z}{1+z_{c}(L_{\gamma})}\right)^{p1}+\left(\frac{1+z}{1+z_{c}(L_{\gamma})}\right)^{p2}\right]^{-1}$ (5)
and
$z_{c}(L_{\gamma})=z_{c}^{*}\cdot(L_{\gamma}/10^{48})^{\alpha}.$ (6)
Here, $z_{c}(L_{\gamma})$ represents the luminosity-dependent redshift at which the evolutionary trend changes sign (from positive to negative), with $z_{c}^{*}$ denoting the redshift peak for an FSRQ of luminosity $10^{48}$ erg s-1 and the parameter $\alpha$ controls how the peak redshift shifts with luminosity. The evolution factor $e(z,L_{\gamma})$ is parametrized by two evolutionary indices, $p_{1}$ and $p_{2}$, which describe the redshift evolution below and above the luminosity-dependent peak redshift $z_{c}(L_{\gamma})$, respectively. The values of all parameters are adopted from Ref. [50]. $\Phi(L_{\gamma})$ adopts the same double power-law form as in Eq. (2). This parametrisation is similar to that proposed by Ref. [43], but it is continuous around the redshift peak $z_{c}(L_{\gamma})$. Such continuity is advantageous for fitting algorithms that use derivatives of the function to locate the minimum.
II.2 BL Lacertae (BL Lac) Objects
In this case, we use the intrinsic distribution of photon indices following Ref. [52], where the photon index $\Gamma$ is assumed to follow a Gaussian distribution for a given redshift $z$ and luminosity $L_{\gamma}$ and hence the luminosity function becomes
$\Phi(L_{\gamma},z,\Gamma)\propto\exp\left[-\frac{(\Gamma-\mu(L_{\gamma}))^{2}}{2\sigma^{2}}\right],$ (7)
where $\mu$ and $\sigma$ are the Gaussian mean and dispersion, respectively. Previous works [53, 54] have suggested that the mean photon index may correlate with luminosity, and thus we allow $\mu$ to vary with $L_{\gamma}$ following the blazar sequence prescription of Ref. [52], given by
$\mu(L_{\gamma})=\mu^{*}+\beta\,[\log_{10}(L_{\gamma})-46],$ (8)
where $\mu^{*}$ is the mean photon index at a reference luminosity $L_{\gamma}=10^{46}\,\mathrm{erg\,s^{-1}}$, and $\beta$ quantifies the slope of the luminosity–spectral index correlation. The LF at redshift $z=0$ is modelled as a smoothly connected double power law, multiplied by the photon index distribution, given in Eq. (7), as
$\Phi(L_{\gamma},z=0,\Gamma)=\frac{A}{2.303\,L_{\gamma}}$ $\left[\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma_{1}}+\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma 2}\right]^{-1}$
$\times\exp\left[-\,\frac{(\Gamma-\mu(L_{\gamma}))^{2}}{2\sigma^{2}}\right].$ (9)
As in the FSRQs case, in this case the relation between the observed luminosity $L_{\gamma}$ and the intrinsic luminosity $L_{\gamma}^{\prime}$ is $L_{\gamma}=\delta_{\rm e}^{4}L_{\gamma}^{\prime}$ [52, 51]. Also, we utilise here the same formalism as performed in Eqs. (4)–(6) and the best fit values of the parameters are adopted from Ref. [52].
III Diffusion of Cosmic Rays in Turbulent Magnetic Fields
Modelling extragalactic magnetic fields remains a complex task due to several uncertainties and limitations [55]. The exact strength and structure of these fields are still not well known and can vary greatly across different parts of extragalactic space [56, 57]. To make the study more manageable, we consider the propagation of CRs in a uniform and turbulent extragalactic magnetic field. Such a field is described by two key parameters: its root mean square (RMS) strength $B$ and its coherence length $l_{\text{c}}$. The RMS strength, defined as $\sqrt{\langle B^{2}(x)\rangle}$, may be as weak as $10^{-16}$ G in cosmic voids [58], increase up to $\gtrsim 10$ nG in galaxy clusters, and typically reach about $1$ nG in filaments [59, 60, 61]. The coherence length $l_{\text{c}}$ generally lies between $0.01$ Mpc and $1$ Mpc [3]. The effective Larmor radius $r_{\text{L}}$ of a charged particle with charge $Ze$ and energy $E$ moving through a turbulent magnetic field of strength $B$ can then be written as
$r_{\text{L}}=\frac{E}{ZeB}\simeq 1.1\,\frac{E/\text{EeV}}{ZB/\text{nG}}\;\text{Mpc}.$ (10)
The idea of critical energy is essential for describing how charged particles diffuse in magnetic fields. It is the energy at which the coherence length of a particle with charge $Ze$ becomes equal to its Larmor radius, i.e., $r_{\text{L}}(E_{\text{c}})=l_{\text{c}}$. Hence, the critical energy is written as
$E_{\text{c}}=ZeBl_{\text{c}}\simeq 0.9Z\,\frac{B}{\text{nG}}\,\frac{l_{\text{c}}}{\text{Mpc}}\;\text{EeV}.$ (11)
This energy separates two different diffusion regimes: resonant diffusion at energies below $E_{\text{c}}$ and non-resonant diffusion at energies above $E_{\text{c}}$.
The energy-dependent diffusion coefficient $D$ is expressed as [62]
$D(E)\simeq\frac{c\,l_{\text{c}}}{3}\left[4\left(\frac{E}{E_{\text{c}}}\right)^{2}+a_{\text{I}}\left(\frac{E}{E_{\text{c}}}\right)+a_{\text{L}}\left(\frac{E}{E_{\text{c}}}\right)^{2-\gamma}\right],$ (12)
where $\gamma$ is the spectral index, and $a_{\text{I}}$ and $a_{\text{L}}$ are constants. For a Kolmogorov-type turbulence spectrum, $\gamma=5/3$ with $a_{\text{L}}\approx 0.23$ and $a_{\text{I}}\approx 0.9$. As noted earlier, the diffusion length $l_{\text{D}}$ represents the typical distance over which a particle’s total deflection becomes roughly one radian, defined as $l_{\text{D}}=3D/c$. Throughout this work, we assume a homogeneous turbulent intergalactic magnetic field with RMS strength $B_{\rm rms}=1~\mathrm{nG}$ and coherence length $l_{\rm c}=0.1~\mathrm{Mpc}$, corresponding to the maximum scale of a Kolmogorov turbulence spectrum.
In the diffusive regime, the transport equation for UHE particles moving through an expanding Universe from a source at position $x_{\text{s}}$ is given by [63]
$\frac{\partial n}{\partial t}+3H(t)\,n-b(E,t)\,\frac{\partial n}{\partial E}$ $-n\,\frac{\partial n}{\partial E}-\frac{D(E,t)}{a^{2}(t)}\,\nabla^{2}n$
$=\frac{\mathcal{N}(E,t)}{a^{3}(t)}\,\delta^{3}({x}-{x}_{\text{s}}),$ (13)
where the Hubble parameter is $H(t)=\dot{a}(t)/a(t)$, with $\dot{a}(t)$ representing the rate of change of the scale factor $a(t)$ over cosmic time $t$. The coordinates x{x} denote comoving positions, $n$ is the particle density, and $\mathcal{N}(E)$ represents the source emissivity. At a time $t$ corresponding to redshift $z$, the separation between the source and the particle is given by $r_{\text{s}}={x}-{x_{\text{s}}}$. Energy losses caused by the expansion of the Universe and interactions with the CMB are included in this formulation as
$\frac{dE}{dt}=-\,b(E,t),\quad b(E,t)=H(t)E+b_{\text{int}}(E).$ (14)
Here, $H(t)E$ corresponds to adiabatic energy losses due to the expansion of the Universe, while $b_{\text{int}}(E)$ represents energy losses arising from interactions. In this work, all interaction losses are calculated using only the CMB as the target photon field. For proton primaries, we include energy losses due to pair production and photopion production upon interactions with the CMB, treated in the continuous energy-loss approximation [62]. For nuclear primaries ($A>1$), we include pair production and photodisintegration caused by interactions with the CMB only; photon fields from the extragalactic background light (EBL) are not included. This approximation is appropriate for the class of sources considered here, which are able to accelerate nuclei to UHEs without strong photodisintegration [64, 65]. According to the Ref. [66], photodisintegration largely preserves the Lorentz factor and rigidity of the main nuclear fragment. As a result, this process does not change the diffusion coefficient or the overall form of particle propagation, and its effect enters only through the nuclear survival probability. Similar treatments have also been discussed in the context of modified and alternative gravity models [67, 68, 19].
The general solution to Eq. (III) was derived in Ref. [63] and is written as
$n(E,r_{\text{s}})=\int_{0}^{z_{i}}dz\,\bigg|\frac{dt}{dz}\bigg|\,\mathcal{N}(E_{\text{g}},z)\,\frac{\textrm{exp}\left[-r_{\text{s}}^{2}/4\lambda^{2}\right]}{(4\pi\lambda^{2})^{3/2}}\,\frac{dE_{\text{g}}}{dE},$ (15)
where $\lambda$ is the Syrovatskii variable, defined as [69]
$\lambda^{2}(E,z)=\int_{0}^{z}dz\left|\frac{dt}{dz}\right|(1+z)^{2}D(E_{\text{g}},z).$ (16)
[…]
$\left|\frac{dt}{dz}\right|=\frac{1}{H_{0}\,(1+z)\,\sqrt{(1+z)^{3}\,\Omega_{m}+\Omega_{\Lambda}}},$ (17)
[…]
$\xi(E,r_{\text{s}})=\frac{4\pi r_{\text{s}}^{2}c\,n(E,r_{\text{s}})}{\mathcal{N}(E)}.$ (18)
IV Flux of UHECRs
The diffusion of CRs in TMFs has been studied extensively by many authors [63, 72, 73, 74, 3, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. Berezinsky and Gazizov [85, 63] generalised the Syrovatskii solution [69] to examine the diffusion of protons in an expanding Universe. The flux from a CR source at a distance $r_{\text{s}}$, much larger than the diffusion length $l_{\text{D}}$, can be obtained by solving the diffusion equation in an expanding Universe [63], leading to the expression [86]:
$J(E)=\frac{c}{4\pi}\int_{0}^{z_{\text{max}}}dz\,$ $\left|\frac{dt}{dz}\right|\,\mathcal{N}\left[E_{\text{g}}(E,z),z\right]$ (19)
$\frac{\exp\left[-r_{\text{s}}^{2}/(4\lambda^{2})\right]}{(4\pi\lambda^{2})^{3/2}}\frac{dE_{\text{g}}}{dE},$
where $z_{\text{max}}$ is the maximum redshift at which the source starts emitting CRs, and $E_{\text{g}}$ is the generation energy at redshift $z$ corresponding to energy $E$ at $z=0$. The total source emissivity $\mathcal{N}$ is obtained by summing the charge-specific emissivities $\mathcal{N}_{\text{Z}}$ for different nuclei. The charge-specific emissivity follows a power-law with a rigidity cutoff $ZE_{\text{max}}$ and is given by [66]
$\mathcal{N}_{\text{Z}}(E,z)=\varepsilon_{\text{Z}}f(z)E^{-\gamma}/\cosh(E/ZE_{\text{max}}),$ (20)
where $\varepsilon_{\text{Z}}$ denotes the relative contribution of nuclei with charge $Z$ to the CR flux, and $f(z)$ describes the evolution of source emissivity with redshift $z$. In our study, we improve the description of $f(z)$ by adopting the LDDE model for FSRQs and BL Lacs derived by Ajello et al. [50, 52]. Instead of using a simple phenomenological form, often taken as a power-law or broken power-law in $(1+z)$, we employ a physically motivated redshift evolution based on the observed FSRQ luminosity function. In this framework, the comoving source density is expressed as
$f(z)=E^{-\gamma}\int_{L_{\min}}^{L_{\max}}\Phi(L,z)\,L\,dL,$ (21)
where $\Phi(L,z)$ is the luminosity function of the sources and $L_{\text{min}}$, $L_{\text{max}}$ are the luminosity limits. This integration properly incorporates the redshift and luminosity dependence of the AGN population. This approach ensures that the source evolution is not arbitrary, but directly informed by observational data, providing a more realistic representation of UHECR source emissivity. In contrast, simple phenomenological forms such as $f(z)\propto(1+z)^{m}$ or broken power-laws do not account for the luminosity dependence of AGNs and may result in oversimplified or biased estimates of source evolution. The Ajello LDDE model thus offers a robust, observation-based framework for modelling UHECR source evolution.
Equation (19) can be generalised to nuclei in terms of their rigidity [66]. For multiple sources, we apply the propagation theorem [87] to sum contributions from all sources, which satisfies
$\int_{0}^{\infty}dr,4\pi r^{2}\frac{\exp\left[-r^{2}/(4\lambda^{2})\right]}{(4\pi\lambda^{2})^{3/2}}=1.$ (22)
To examine the effect of finite source distances on flux suppression, we compute the sum over a given set of distance distributions. These assume a uniform source density, with distances from the observer given by [66, 86, 88]
$r_{\text{i}}=\left(\frac{3}{4\pi}\right)^{1/3}\,d_{\text{s}}\,\frac{\Gamma(i+1/3)}{(i-1)!},$ (23)
where $d_{\text{s}}$ is the separation distance between sources. For a discrete source distribution, summing over all sources yields a suppression factor [66, 86]
$F\equiv\frac{1}{n_{\text{s}}}\sum_{i}\frac{\exp\left[-r_{\text{i}}^{2}/4\lambda^{2}\right]}{(4\pi\lambda^{2})^{3/2}},$ (24)
in place of Eq. (22), where $n_{\text{s}}$ is the source density. Using Eq. (19), after summing over all sources, the modified flux for an ensemble of sources can then be expressed as
$J_{\text{mod}}(E)\simeq\frac{R_{\text{H}}n_{\text{s}}}{4\pi}\!$ $\int_{0}^{z_{\text{max}}}\!\!\!\!\!dz\,(1+z)^{-1}\Big|\frac{dt}{dz}\Big|$
$\times\mathcal{N}\left[E_{\text{g}}(E,z),z\right]\,\frac{dE_{\text{g}}}{dE}\,F,$ (25)
where $R_{\text{H}}=c/H_{0}=4.3$ Gpc is known as the Hubble radius. We can rewrite Eq. (16) in terms of $R_{\text{H}}$ and from Eq. (12) as
$\lambda^{2}(E,z)=$ $\,\frac{H_{0}R_{\text{H}}l_{\text{c}}}{3}\int_{0}^{z}\!\!dz\,\bigg|\frac{dt}{dz}\bigg|(1+z)^{2}\bigg[4\left(\frac{(1+z)\,E}{E_{\text{c}}}\right)^{2}$
$+a_{\text{I}}\left(\frac{(1+z)\,E}{E_{\text{c}}}\right)+a_{\text{L}}\left(\frac{(1+z)\,E}{E_{\text{c}}}\right)^{2-\gamma}\bigg].$ (26)
Figure 1: Fluxes of UHECRs and $X_{\text{max}}$ values for a mixed composition scenario from FSRQs (top panel) and BL Lac objects (bottom panel), obtained by using different source separation distances (see text) and varying relative abundances of nuclear species. The observational values of flux of these two experiments are taken from [89] and those for $X_{\text{max}}$ are from [90, 91].
For the mixed composition case, we draw Fig. 1 for the fluxes along with the mean depths of shower maximum $\langle X_{\text{max}}\rangle$ as predicted by our theoretical calculations. The nuclear abundance of p, He, N, Si and Fe is mentioned at the top of the plots. Since we use two different AGN source population, so we get two different separation distance in fitting the numerical results with Auger and TA data [89]. The corresponding values of the source separation distances are $160$ kpc and $36$ Mpc for FSRQs and BL Lac objects, respectively. The depth of the shower maximum, $X_{\text{max}}$, is determined using a parametrization based on air shower physics [92, 93]. For a nucleus with mass number $A$ and energy $E$, it is expressed as [92, 93].
$X_{\text{max}}(E,A)=X_{0}+\nu~\text{log}_{10}\left(\frac{E}{A}\right),$ (27)
where $X_{0}$ and $\nu$ are parameters influenced by hadronic interactions [92]. In our analysis, we have used the values as $X_{0}=700\,\text{g cm}^{-2}$ and $\nu=50\,\text{g cm}^{-2}$ [93, 94].
Refer to captionRefer to caption
Figure 2: Mollweide skymaps of the integral UHECRs flux in the range of $10\leq E\leq 100~\mathrm{EeV}$ for synthetic FSRQ (left) and BL Lac (right) populations. The maps assume a uniform comoving distribution of hypothetical sources, with no observational catalogues used. A $2^{\circ}$ Gaussian smoothing is applied. The colour scale shows $E^{3}J(E)$ in $\mathrm{eV^{2}\,m^{-2}\,s^{-1}\,sr^{-1}}$.
For scenarios involving a mixed composition of CRs, the flux-weighted mean depth of the shower maximum is computed as
$\langle X_{\text{max}}\rangle=\frac{\sum\limits_{i}J_{i}(E)\cdot X_{\text{max, i}}(E,A_{i})}{\sum\limits_{i}J_{i}(E)}.$ (28)
Here, $J_{i}(E)$ denotes the flux of each nuclear species.
We provide the $\chi^{2}$ test for both source models with Auger and TA data, and it is defined as
$\chi^{2}=\sum_{i}\frac{(\text{J}^{i}_{\text{th}}-\text{J}^{i}_{\text{obs}})^{2}}{\sigma_{i}^{2}},$ (29)
where $\text{J}^{i}_{\text{th}}$ are the theoretical values of flux obtained from our calculations and $\text{J}^{i}_{\text{obs}}$ are the observed flux values, which are obtained from the Auger and the TA experiment. $\sigma$ denotes the total uncertainty in each energy bin, incorporating both statistical and systematic errors. The corresponding reduced $\chi^{2}$ values with the combined Auger and TA data are $2.41$ and $2.63$ for FSRQs and BL Lac objects respectively.
Fig. 1 also displays the fluxes for a mixed composition of nuclei along with the corresponding $\langle X_{\text{max}}\rangles$ and $\sigma\left(X_{\text{max}}\right)$s at the redshift $z_{\text{max}}=1$. The parametrisations show deviations from a purely proton composition, and the influence of the astrophysical source model is also apparent. The observational data for $\langle X_{\text{max}}\rangles$ from Auger and TA are taken from Refs. [90, 91]. These datasets provide $\langle X_{\text{max}}\rangle$ values along with statistical and systematic errors. The red and blue bands in $\langle X_{\text{max}}\rangles$ and $\sigma\left(X_{\text{max}}\right)$s plots denote the systematic errors of Auger and TA data, respectively. In the $\langle X_{\text{max}}\rangle$ and $\sigma\left(X_{\text{max}}\right)$ plots, the statistical uncertainties are shown as shaded bands. The good agreement between our results and the observational data supports the validity of the source models used in reproducing the CR spectra. For readability, we include Table 1, which summarizes the key parameters used in fitting the UHECR energy spectrum with the Auger and TA observational data for the FSRQ and BL Lac source classes.
It should be mentioned that the absolute $\langle X_{\max}\rangle$ values measured by the Auger and the TA should not be compared directly. As demonstrated in joint Auger-TA analyses, differences in detector design, event-selection criteria, and reconstruction methods introduce experiment-dependent offsets in $\langle X_{\max}\rangle$ that are unrelated to mass composition. Consequently, composition inference is based on the behaviour of $\langle X_{\max}\rangle$ and $\sigma(X_{\max})$ relative to single-species expectations and on the energy dependence of these observables, rather than on a direct comparison between these two datasets.
Table 1: Comparison of parameters for FSRQs and BL Lac objects. $\varepsilon$ denotes the relative contribution from the nuclei.
Parameters FSRQ BL Lac
$d_{\text{s}}$ $160$ kpc $36$ Mpc
$E_{\text{max}}$ (EeV) $5Z$ $5Z$
$\chi^{2}_{\text{red}}$ $2.41$ $2.63$
d.o.f. 1919 $16$
$\varepsilon_{\text{p}}$ $0.5$ $0.45$
$\varepsilon_{\text{He}}$ $0.25$ $0.28$
$\varepsilon_{\text{N}}$ $0.15$ $0.15$
$\varepsilon_{\text{Si}}$ $0.06$ $0.07$
$\varepsilon_{\text{Fe}}$ $0.04$ $0.05$
To describe the angular distribution of the UHECRs flux, we generate full-sky intensity maps using the HEALPix framework [95] with a resolution parameter of ${\tt Nside}=256$, corresponding to $N_{\text{pix}}=12\times N_{\text{side}}^{2}$ pixels covering the full sky. In this work, all skymaps are built using a uniform comoving distribution of synthetic sources representing the FSRQ and BL Lac populations. No observational source catalogues are used. The positions of the simulated sources are generated randomly over the entire sky in right ascension ($0^{\circ}-360^{\circ}$) and declination ($90^{\circ}-+90^{\circ}$), and then converted to the corresponding HEALPix coordinates $(\theta,\phi)$. For each simulated source, we calculate its contribution to the integral UHECRs flux within the energy range $10\leq E\leq 100~\mathrm{EeV}$ shown in the maps. This is done by numerically integrating the propagated energy spectrum from the chosen energy threshold up to the maximum injection energy, and the resulting flux is added to the appropriate HEALPix pixel. Magnetic deflections are included through the diffusion coefficient $D(E)$, which depends on the strength and coherence length of the intergalactic magnetic field. These parameters in the Syrovatskii variable $\lambda^{2}(E,z)$ describe how the flux from each source is reduced with distance due to magnetic spreading during propagation. To reduce pixel-to-pixel numerical noise and to make large-scale anisotropy patterns easier to see, the final maps are smoothed using a Gaussian function with a width of $2^{\circ}$.
This construction is meant to be illustrative and is not intended to reproduce the observed UHECRs sky. Since the source population is uniformly distributed and entirely synthetic, any localized regions of higher flux (“hotspots”) arise naturally from statistical effects. In the energy range $10\leq E\leq 100~\mathrm{EeV}$, the propagation horizon is limited, so only a small number of nearby synthetic sources contribute significantly to the total flux. This produces apparent hotspots in the skymaps. These features, therefore, reflect the geometry of the simulated source distribution rather than real astrophysical anisotropies. Fig. 2 shows the resulting full-sky integral flux maps for the FSRQ and BL Lac scenarios. The FSRQ case exhibits stronger anisotropy, which is consistent with its higher typical luminosities and LDDE evolution, while the BL Lac case leads to a smoother flux distribution. The hotspots seen in the maps are therefore a natural outcome of the synthetic source model.
Table 2: Nearby galaxies and extragalactic sources within $1^{\circ}$ of the hotspot located at (RA = 253.12, Dec = $87.08^{\circ}$) for FSRQ case.
Object Name Type Redshift ($z$) Separation () RA () Dec ()
ESO 1-5 G/AG? $0.016732$ 0.963 133.27 88.09-88.09
HD 89592 RG* $2.20\times 10^{-5}$ 0.503 146.63 88.36-88.36
HD 99357 RG* $9.69\times 10^{-5}$ 0.703 166.62 88.96-88.96
TYC 9518-226-1 RG* $1.11\times 10^{-4}$ 0.887 147.23 89.07-89.07
CPD-87 200 RG* $1.10\times 10^{-4}$ 0.832 189.49 88.11-88.11
CPD-87 210 RG* $8.04\times 10^{-5}$ 0.913 194.36 88.33-88.33
TYC 9514-397-1 RG* $1.37\times 10^{-5}$ 0.508 148.30 88.09-88.09
TYC 9515-917-1 RG* $8.66\times 10^{-5}$ 0.561 156.23 87.76-87.76
TYC 9515-1109-1 RG* $1.05\times 10^{-4}$ 0.859 163.49 87.40-87.40
TYC 9516-329-1 RG* $2.04\times 10^{-5}$ 0.891 184.90 87.76-87.76
TYC 9518-181-1 RG* $1.18\times 10^{-4}$ 0.963 130.54 88.72-88.72
UCAC4 014-002450 RG* $1.45\times 10^{-4}$ 0.916 156.23 87.38-87.38
UCAC4 013-002324 RG* $2.20\times 10^{-4}$ 0.904 155.85 87.40-87.40
UCAC4 011-002319 RG* $1.38\times 10^{-4}$ 0.461 173.47 87.94-87.94
Table 3: Nearby galaxies and extragalactic sources within $1^{\circ}$ of hotspot located at (RA = 286.88, Dec = 87.08) for BL Lac case.
Object Name Type Redshift ($z$) Separation () RA () Dec ()
2MASX J15433127-8740427 G $0.04313$ 0.771 235.88 87.68-87.68
ESO 1-6 G $0.007538$ 0.909 216.83 87.77-87.77
ESO 1-7 G $0.016581$ 0.650 255.69 88.92-88.92
NOGG P1 815 GrG $0.006525$ 0.917 224.90 87.61-87.61
[TSK2008] 1026 GrG $0.007574$ 0.898 225.41 87.62-87.62
CPD-88 126 RG* $2.77\times 10^{-5}$ 0.551 227.16 88.97-88.97
TYC 9520-151-1 RG* $1.91\times 10^{-5}$ 0.968 200.43 88.36-88.36
TYC 9517-1108-1 RG* $1.19\times 10^{-5}$ 0.421 233.68 88.03-88.03
TYC 9517-1326-1 RG* $4.50\times 10^{-5}$ 0.934 218.49 87.70-87.70
TYC 9517-1406-1 RG* $3.70\times 10^{-6}$ 0.652 224.20 87.91-87.91
TYC 9521-317-1 RG* $2.24\times 10^{-5}$ 0.599 212.90 88.60-88.60
TYC 9530-981-1 RG* $1.77\times 10^{-4}$ 0.568 241.42 87.91-87.91
To identify galaxies or AGNs in the vicinity of the detected hotspots, we performed a positional cross-matching using the astroquery.simbad module. The coordinates of each hotspot (RA and Dec) were used as input to query the SIMBAD astronomical database within a circular search region of radius $1^{\circ}$. Since the skymaps are produced entirely from synthetic sources that are uniformly distributed and therefore do not contain any observational information. Hence, any comparison with the positions of known AGNs from catalogues such as SIMBAD is made only for reference. This comparison is used simply to show which well-known extragalactic objects happen to lie close to the directions of enhanced flux in the model skymaps. However, the visual overlap between hotspots and catalogue objects indicates the efficiency of the predictability of the model considered here.
A custom SIMBAD query was configured to retrieve essential information such as the main identifier (main_id), object type (otypes), and equatorial coordinates (RA, Dec). The returned catalogue entries were then filtered to include only extragalactic sources, specifically those classified as galaxies or active nuclei. The following SIMBAD object types were considered: G (Galaxy), AGN, Radio (Radio Source), Sy (Seyfert Galaxy), and LINER (Low-Ionization Nuclear Emission-line Region). For each candidate object, the angular separation from the hotspot position was calculated using the astropy.coordinates.SkyCoord module, and only those lying within $1^{\circ}$ were considered. This method ensures the identification of nearby galaxies or AGN-like sources that could be potential counterparts or contributors to the observed hotspot regions in the sky. Numerous entries are classified as RG* (Red Giant Stars), and some are AG? (possible active galaxy) and GrG (galaxy group).
For FSRQs, using the above method, we searched for extragalactic sources within an angular radius of $1^{\circ}$ around the hotspot located at $(\mathrm{RA}=253.12^{\circ},\,\mathrm{Dec}=-87.08^{\circ})$. Notable entries from ESO 1-5, CPD, TYC, and UCAC are found in this region. The presence of nearby galaxies in the vicinity of this hotspot may suggest possible associations between these extragalactic sources and the observed high-energy CRs flux enhancement in this direction. For the BL Lac objects, among the identified sources within $1^{\circ}$ of the hotspot located at RA $286.88^{\circ}$, Dec $87.08^{\circ}$, several prominent extragalactic objects were found. A few possible active galaxies (AG?) like ESO 1-6 and ESO 1-7 were identified, suggesting the presence of AGNs in the vicinity. The catalogue also includes several RG* from CPD and TYC catalogue. Overall, the spatial correlation indicates that the region is populated with a mix of normal and active galaxies, as well as nearby stellar contaminants. For both FSRQs and BL Lac objects, these nearby galaxies or extragalactic sources are given in Tables 2 and 3.
V Summary and Conclusion
[…]
We constructed full-sky maps of the integral UHECR flux in the energy range $10\leq E\leq 100~\mathrm{EeV}$ using the HEALPix framework with an angular resolution of ${\tt Nside}=256$. The maps were generated assuming a uniform comoving distribution of synthetic sources for each source class, without using any observational source catalogue. Source positions were assigned randomly over the full sky, and each source contribution was added to the corresponding HEALPix pixel. The integral flux above the chosen energy threshold was obtained by numerically integrating the propagated spectrum up to the maximum injection energy. To suppress small-scale numerical fluctuations and highlight large-scale anisotropy, the maps were smoothed with a Gaussian function with a width of $2^{\circ}$. These skymaps are intended for illustration only, as the underlying source distribution is entirely synthetic. At energies above $\sim 10~\mathrm{EeV}$, the limited propagation horizon implies that only a few nearby sources dominate the flux, naturally producing localized excesses or “hotspots”. Such features arise from proximity effects in the simulations and do not represent real astrophysical anisotropies.
To provide contextual information on the sky regions corresponding to the synthetic hotspots, we performed a positional cross-matching with the SIMBAD astronomical database using the astroquery.simbad module. This analysis was carried out solely as an exploratory and orientational exercise. We emphasise that the skymaps themselves contain no observational directional information, and the subsequent comparison with catalogued galaxies or active galactic nuclei does not imply any physical association or source identification. The cross-matching serves only to illustrate which known extragalactic objects happen to lie near the directions of the model-generated flux enhancements. The identified nearby galaxies and AGN-like sources are summarised in Tables 2 and 3.
Overall, our results demonstrate how energy-dependent magnetic propagation effects and finite CRs horizons can naturally produce anisotropy like features in UHECR skymaps, even for uniformly distributed source populations. The framework developed here provides a self–consistent basis for interpreting such features in synthetic models and for exploring how different source classes may contribute to the observed UHECR sky.
Acknowledgement
UDG is thankful to the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, India for the Visiting Associateship of the institute.